# Gen-C: Populating Virtual Worlds with Generative Crowds

### v1

@@ 4.3 Crowd Scenario Graph Representation
Figure 2: Structure of a crowd scenario graph. Each node contains an action label, a location category, an agent ID, and timestep as $\alpha_{i}$_$t$.
@@ 5 Learning crowd scenario graphs
While a synthetic crowd scenario graph can be used to drive a crowd of simulated agents, ideally we want to be able to generalize without having to manually compute a new graph from scratch every time the input conditions change. To do so, given a dataset consisting of crowd scenario graphs, we will employ a generative model that can reason about the graphs and generate new ones.
@@ 5.1 Data Preprocessing
Before training, we apply some preprocessing steps to prepare our data. For each sample $(S_{in},SG_{i},agents_{i})$, we treat subgraph $SG_{i}$ as an individual undirected graph $G_{i}=(V_{i},E_{i})$, where $n=|V_{i}|$ (number of nodes) and $m=|E_{i}|$ (number of edges). The adjacency matrix $\mathcal{A}_{G_{i}}\in\mathbb{R}^{n\times n}$ of a graph $G_{i}$ is a symmetric matrix that encodes edge information in the graph. The value of the $i^{th}$ row and $j^{th}$ column is equal to 00 if there is no edge between $v_{i}$ and $v_{j}$. Otherwise, it set to $1$ or $-1$, if edge’s type is “sequence” or “share” respectively.
@@
Node Features. We annotate the graphs nodes using feature vectors. The feature vector of each node includes: the action, location, unique agent id, sequence index and interaction status. First, we use action and location as two one-hot vectors (sizes 15 and 8) from the available actions $\mathcal{A}ct$ and available location categories $\mathcal{L}oc$. For agent ID (${\alpha_{i}}$) and timestep $t$ in sequence, we set the maximum number of agents per subgraph equal to 6 and the maximum sequence length to 10; these choices reduce the complexity of generated graphs.
We construct another two one-hot encodings (sizes 6 and 10) for them too. We encode the interaction status of the current node, indicating whether the current action is shared or not with another one-hot vector (size 2). To generate the final node features, we use embedding layers to ensure that node features are dense and learnable representations suitable for downstream tasks. Each encoding is then passed through its respective embedding layer: actions ($\mathbb{R}^{15\times 6}$), locations ($\mathbb{R}^{8\times 4}$), agents ($\mathbb{R}^{6\times 5}$), timestep ($\mathbb{R}^{10\times 4}$), and interaction status ($\mathbb{R}^{2\times 2}$). The embeddings are concatenated to form the final node feature representation $\mathbf{X}\in\mathbb{R}^{n\times d}$ where $d$ is the feature dimension size. Additionally, we include the top 4 eigenvectors of the normalized graph Laplacian as node features to encode global structural properties [60]. Thus, the final node feature dimension becomes $d=21+4=25$.
Textual Condition. We convert $S_{in}$ to textual condition by first detecting and emphasizing verbs in the text to highlight action-relevant information; we utilize the spaCy library [61]. Each $S_{in}$ is augmented by prefixing it with a list of extracted verbs, and its embedding is generated using the pretrained Sentence Transformer model all-MiniLM-L6-v2 [62].
@@ 5.2 Training
Graph Structure Decoder. Reconstructs the padded graph adjacency matrix $\mathcal{A}_{G^{\prime}}$ from $Z_{S}$; we set the maximum number of nodes equal to 40. It uses a MLP to transform the concatenated latent and conditional representations into a flattened adjacency matrix. An MLP maps the concatenated latent and conditional representations to the upper triangular part of the adjacency matrix, which is symmetrized to produce the full matrix.
Features Decoder. Reconstructs node features $\mathbf{X}_{G^{\prime}}$ from $Z_{F}$, and $\mathcal{A}_{G}$. The decoder employs a series of GINE layers, where the input to each layer is a combination of the expanded $Z_{F}$, $\mathbf{C}$. Each GINEConv layer applies a message-passing operation, followed by non-linear transformations using LeakyReLU activations and LayerNorm for stability. The final layer projects the hidden representations into the node feature space, with size 23.
Condition Net. Both VGAE models use the same condition network structure that encodes the embeddings of text ($S_{in}$) and number of agents to produce the condition vector $C$. We utilize $\mathbf{C}$ to train two Prior Networks that parameterize the prior distributions $p(Z|C)$ of the latent variables of VGAE-S and VGAE-F, computing their respective means $\mu_{\text{prior}}$ and standard deviations $\sigma_{\text{prior}}$. Additionally, $C$ is integrated into both the encoding and decoding phases. In the encoder, it is concatenated with the graph-level features embedding to produce the posterior distribution $q(Z|X,C)$, parameterized by $\mu$ and $\sigma$., while during decoding it guides the reconstruction by conditioning the latent space.
@@
where SmoothL1 denotes a smooth L1-error and CELoss denotes the cross-entropy loss. The KL divergence loss $\mathcal{L}_{KL}$ measures the difference between the posterior distribution $q(Z|X,C)$ and the prior distribution $p(Z|C)$, and is computed by:
@@
[…]
@@ 6 Text-Conditioned Scenario Generation
We parse the generated graphs to collect a sequence of actions-locations and interactions per agent. Then, to generate a simulation, a simulation environment can be either (a) manually provided by the user or (b) generated with a simple LLM query similar to $Q2$, as described in Section 4.1. Given the environment, we run simulations of the generated scenarios in Unity [5]. We provide visual results in Fig. 7, while related animations are also presented in the supplementary video.
@@ 7 Experiments and Evaluation
We train our model on two synthetic datasets representing crowd scenarios for a University Campus and a Train Station. For each dataset, we begin by defining 100 unique input sentences describing crowd scenarios; an example input is “Outside the station, passengers sit on benches to read a newspaper and talk on the phone.” Using the process detailed in Section 4.1, we generate 5 variations for each sentence, resulting in 500 crowd scenarios per dataset. Graph datasets are then constructed as outlined in Section 4.2 and Section 4.3, with the University Campus dataset containing $135k$ subgraphs and the Train Station dataset $110k$. The training data and models will be released upon acceptance. We split each dataset into 75% for training, 15% for validation, and 10% for testing, and train a separate model per dataset. For model training (cf. Fig. 4), we use a graph encoder consisted of 2 GINE layers with a hidden dimension of 96, each followed by a dropout layer with a rate of .2. Both VGAE-S and VGAE-F have a latent dimension of 16. The graph structure decoder uses 3 linear layers with a hidden dimension of 128, applying a tanh activation after the final layer. The feature decoder employs 3 GINE layers with a hidden dimension of 128, followed by a layer mapping to a 23-dimensional feature space. The condition network projects text embeddings from 384 to 128 dimensions and agent embeddings from 32 to 128, combining them into a conditional vector of size 128. The prior network maps this conditional vector to 16 dimensions using 2 linear layers. LeakyReLU activations are used throughout, with BatchNorm in the encoder and LayerNorm in both decoders for stability. The models are trained for 2,000 epochs with a batch size of 256, a learning rate of 0.001, and the Adam optimizer with a weight decay of 0.0004. Every cycle for the $\beta$ annealing lasts 200 epochs. Training each model required approximately 4 hours on a single NVIDIA RTX 4070 Ti GPU.
Quantitative Results.
@@
We also evaluate the diversity of feature generation in our model. Using the same textual description $S_{in}$, we sample 1k variables from the latent space of VGAE-F and collect each label’s frequency. The results for four user-input scenarios are presented in Fig. 6, demonstrating that the model produces diverse outputs rather than being limited to a fixed set. Moreover, the generated results remain consistent with the input condition $S_{in}$.
@@
Degree .2578 .0053 .0045 .4710 .0045 .0055
Clustering Coeff. .5143 .0203 .0046 .4511 .0096 .0074
Diameter .1996 .1425 .1200 .2642 .0369 .0820
Avg. Path Length .1877 .2244 .0360 .6330 .1005 .0500
Action .4149 .3009 .0640 .2699 .3078 .1100
Location .0925 .0316 .0027 .0270 .0480 .0088
@@ 8 Discussion and Future Work
We propose Gen-C, a framework designed to populate virtual environments with multiple agents exhibiting diverse high-level actions and interactions, conditioned on textual input. Our approach leverages an LLM and two learned latent spaces to enable the generation of crowd scenario graphs that drive agent behaviors.
Overall, our system generate plausible simulations, allowing the automatic creation of virtual environments with agents exhibiting diverse, high-level actions beyond typical low-level crowd behaviors like goal-seeking and collision avoidance. However, several limitations exist. The simulator used to expand LLM-generated scenarios relies on probabilistic action sequences, thus the sequence of actions may not be optimal. Agents cannot switch actions mid-execution, and action durations are manually defined. Additionally, behaviors are constrained to a predefined list of high-level actions, requiring retraining to accommodate new actions.
For future work, we aim to integrate our framework with a crowd simulation framework, allowing high-level behaviors to control low-level navigation policies, creating more dynamic and realistic simulations. Additionally, we plan to explore combining latent spaces from different datasets, such as "indoor" and "outdoor" scenarios, using techniques such as distillation to enable lifelong learning and dynamic simulation adaptability. We envision that these will allow us to populate virtual worlds with diverse human-like characters, paving the road towards automatic generation of digital crowds that can be applied to a variety of important areas.

### v2

@@ 4.2 Crowd Scenario Graph Representation
Figure 2: Structure of a crowd scenario graph. Each node contains an action label, a location category, an agent ID, and timestep as $\alpha_{i}$_tt.
@@ 5 Learning Crowd Scenario Graphs
While a synthetic LLM-generated crowd scenario graph can be used to drive a crowd of simulated agents, ideally we want to be able to generalize without having to manually query an LLM to compute a new graph from scratch every time the input conditions change. To do so, given a dataset consisting of crowd scenario graphs, we will train a generative model to generated novel graphs.
@@ 5.1 Data Preprocessing
Before training, we apply some preprocessing steps to prepare our data. For each sample $(S_{in},SG_{i},agents_{i})$, we treat subgraph $SG_{i}$ as an individual undirected graph $G_{i}=(V_{i},E_{i})$, where $n=|V_{i}|$ (number of nodes) and $m=|E_{i}|$ (number of edges). The adjacency matrix $\mathcal{A}_{G_{i}}\in\mathbb{R}^{n\times n}$ of a graph $G_{i}$ is a symmetric matrix that encodes edge information in the graph. The value of the $i^{th}$ row and $j^{th}$ column is equal to 0 if there is no edge between $v_{i}$ and $v_{j}$. Otherwise, it set to $1$ or 1-1, if edge’s type is “sequence” or “share” respectively.
@@
Node Features. We annotate the graph nodes using feature vectors. The feature vector of each node includes: the action, location, unique agent id, sequence index and interaction status. First, we use action and location as two one-hot vectors (sizes 13 and 8) from the available actions $\mathcal{A}ct$ and available location categories $\mathcal{L}oc$. For agent ID ($\alpha_{i}$}) and timestep $t$ in sequence, we set the maximum number of agents per subgraph equal to 50 and the maximum sequence length to 45; these choices reduce the complexity of generated graphs. We construct another two one-hot encodings (sizes 6 and 10) for them too. We encode the interaction status of the current node, indicating whether the current action is shared or not with another one-hot vector (size 2). To generate the final node features, we use embedding layers to ensure that node features are dense and learnable representations suitable for downstream tasks. Each encoding is then passed through its respective embedding layer: actions ($\mathbb{R}^{13\times 5}$), locations ($\mathbb{R}^{8\times 4}$), agents ($\mathbb{R}^{6\times 8}$), timestep ($\mathbb{R}^{4}$ using sinusoidal encoding), and interaction status ($\mathbb{R}^{2\times 2}$). The embeddings are concatenated to form the final node feature representation $\mathbf{X}\in\mathbb{R}^{n\times d}$ where $d$ is the feature dimension size. Additionally, we include the top 4 eigenvectors of the normalized graph Laplacian as node features to encode global structural properties [61]. Thus, the final node feature dimension becomes $d=25$.
Textual Condition. We convert $S_{in}$ into a textual condition by computing its embedding using the pretrained Sentence Transformer model all-MiniLM-L6-v2 [62].
@@ 5.2 Training
Graph Structure Decoder. Reconstructs the padded graph adjacency matrix $\mathcal{A}_{G^{\prime}}$ from $Z_{S}$; we set the maximum number of nodes equal to 50. It uses a MLP to transform the concatenated latent and conditional representations into a flattened adjacency matrix. An MLP maps the concatenated latent and conditional representations to the upper triangular part of the adjacency matrix, which is symmetrized to produce the full matrix.
Features Decoder. Reconstructs node features $\mathbf{X}_{G^{\prime}}$ from $Z_{F}$, and $\mathcal{A}_{G}$. The decoder employs a series of GINE layers, where the input to each layer is a combination of the expanded $Z_{F}$, $\mathbf{C}$. Each GINEConv layer applies a message-passing operation, followed by non-linear transformations using LeakyReLU activations and LayerNorm for stability. The final layer projects the hidden representations into the node feature space, with size 21.
Condition Net. Both VGAE models use the same condition network structure that encodes the embeddings of text ($S_{in}$), number of agents, and number of nodes to produce the condition vector $C$. We utilize $\mathbf{C}$ to train two Prior Networks that parameterize the prior distributions $p(Z|C)$ of the latent variables of VGAE-S and VGAE-F, computing their respective means $\mu_{\text{prior}}$ and standard deviations $\sigma_{\text{prior}}$. Additionally, $C$ is integrated into both the encoding and decoding phases. In the encoder, it is concatenated with the graph-level features embedding to produce the posterior distribution $q(Z|X,C)$, parameterized by $\mu$ and $\sigma$., while during decoding it guides the reconstruction by conditioning the latent space.
@@
where SmoothL1 denotes a smooth L1-error and CELoss denotes the cross-entropy loss.
The KL divergence loss $\mathcal{L}_{KL}$ measures the difference between the posterior distribution $q(Z|X,C)$ and the prior distribution $p(Z|C)$, and is computed by:
@@
We use a scale factor of $\beta\in[0,4]$ to balance reconstruction losses $\mathcal{L}_{S}$ and $\mathcal{L}_{F}$ against $\mathcal{L}_{KL}$ and employ a cyclical annealing schedule [64], periodically ramping $\beta$ from 0 its maximum. The final training loss for the two VGAE models are: $\mathcal{L}_{VGAE-G}=\mathcal{L}_{S}+\beta\mathcal{L}_{KL_{S}}$, and $\mathcal{L}_{VGAE-F}=\mathcal{L}_{F}+\beta\mathcal{L}_{KL_{F}}$.
@@ 6 Text-Conditioned Scenario Generation
We parse the generated graphs to collect a sequence of actions-locations and interactions per agent. Then, to generate a simulation, a simulation environment can be either (a) manually provided by the user or (b) generated with a simple LLM query similar to $Q1$, as described in Section 4.1. Given the environment, we run simulations of the generated scenarios in Unity [5]. We provide visual results in Fig. 7, while related animations are also presented in the supplementary video.
@@ 7 Experiments and Evaluation
We train our model on two synthetic datasets representing crowd scenarios for a University Campus and a Train Station. For each dataset, we begin by defining 100 unique input sentences describing crowd scenarios; an example input is “Outside the station, passengers sit on benches to read a newspaper and talk on the phone.” Using the process detailed in Section 4.1, we generate 5 variations for each sentence, resulting in 500 crowd scenarios per dataset. Graph datasets are then constructed as outlined in Section 4.2, with the University Campus dataset containing $12.5k$ subgraphs and the Train Station dataset $13.8k$. The training data and models will be released upon acceptance. We split each dataset into 75% for training, 15% for validation, and 10% for testing, and train a separate model per dataset. For model training (cf. Fig. 4), we use a graph encoder consisted of 2 GINE layers with a hidden dimension of 96, each followed by a dropout layer with a rate of .2. Both VGAE-S and VGAE-F have a latent dimension of 16. The graph structure decoder uses 3 linear layers with a hidden dimension of 128, applying a tanh activation after the final layer. The feature decoder employs 3 GINE layers with a hidden dimension of 128, followed by a layer mapping to a 21-dimensional feature space. The condition network projects text embeddings from 384 to 128 dimensions and agent embeddings from 32 to 128, combining them into a conditional vector of size 128. The prior network maps this conditional vector to 16 dimensions using 2 linear layers. LeakyReLU activations are used throughout, with BatchNorm in the encoder and LayerNorm in both decoders for stability. The models are trained for 2,000 epochs with a batch size of 256, a learning rate of 0.001, and the Adam optimizer with a weight decay of 0.0004. Every cycle for the $\beta$ annealing lasts 200 epochs. Training each model required approximately 4 hours on a single NVIDIA RTX 4070 Ti GPU.
Reconstruction Performance.
@@
Real Data Consistency.
Gen-C uses the knowledge that LLMs acquired from text data to synthesize crowd scenario graphs, reducing reliance on real-world video data. To evaluate this approach, we annotated high-level actions in the Students03 video from the UCY dataset [37], which features 434 students over 3 minutes and 36 seconds. We segmented the video into 10-second windows and recorded the following action frequencies: discuss = 287, meet = 61, talk to phone = 14, carry = 15, wait = 13, sit = 52, wander = 3, and wave at = 5. As not all available actions are present in the real video, actions with zero frequency (e.g., read, look at, queue, enter/exit) are omitted. As shown in the top right of Fig. 6, our model’s action distribution for the University Campus closely aligns with this real data, showing similarly high values for discuss and meet, and low for talk to phone and wait. The model overproduces sit and wander, likely due to the real scene capturing a small-medium size area lacking seating areas.
Action Sequence Analysis.
To analyze the sequence of actions in our data, we perform the following procedure. For each scenario, we gather all graphs from the dataset and count the occurrences of each subsequent action for every current action. The results are shown in Fig. 5. We observe that action transitions are generally intuitive—for example, "discuss" frequently follows "meet", and "wander" often follows "enter/exit", which aligns with behavior when entering a new area. Additionally, certain actions are more prominent based on the scenario: for instance, "carry" shows higher frequencies in the Train Station, as people often carry luggage in this setting.
Figure 5: Action Sequence Tables. We present the frequency of each future action for each action label.
Latent Space Analysis.
To quantify how closely our generated graphs inhabit the same learned latent manifold as the training data, we sampled a batch of graphs from the learned prior and encoded both trained and predicted graphs to obtain two sets of embeddings, $z_{\mathrm{train}}$ and $z_{\mathrm{pred}}$. We visualized these embeddings in Figure 8, where we observe well-mixed clusters with no clear signs of mode collapse. To measure the distributional gap between train and generated embeddings, we computed the Fréchet Distance (FD) [66] between their Gaussian distributions:
$\mathrm{FD}=\|\mu_{\mathrm{train}}-\mu_{\mathrm{pred}}\|_{2}^{2}+\mathrm{Tr}\left(\Sigma_{\mathrm{train}}+\Sigma_{\mathrm{pred}}-2\sqrt{\Sigma_{\mathrm{train}}\Sigma_{\mathrm{pred}}}\right).$
A lower FD indicates a closer alignment between distributions. For feature embeddings, we observe very low FD values (Train Platform: 3.6; University Campus: 6.0), suggesting near-perfect feature recovery. In contrast, the FD for graph structure embeddings is higher (Train Platform: 33.2; University Campus: 19.4), yet still reflects reasonable diversity without significant drift from the real-data manifold.
@@
Degree 1.087 .620 .590 1.120 .107 .110
Clustering Coeff. 2.289 .015 .004 2.170 .0083 .003
Diameter 1.326 .723 .370 1.401 .494 .330
Avg. Path Length 1.602 .310 .150 1.632 .169 .200
Action 0.210 .141 .130 .118 .213 .066
Location 0.059 .032 .019 .021 .0512 .015
@@ 8 Discussion and Future Work
We propose Gen-C, a framework designed to populate virtual environments with multiple agents exhibiting diverse high-level actions and interactions, conditioned on textual input. Our approach leverages an LLM and two learned latent spaces to enable the generation of crowd scenario graphs that drive agent behaviors. Overall, our system generates plausible simulations, allowing the automatic creation of virtual environments with agents exhibiting diverse, high-level actions beyond typical low-level crowd behaviors like goal-seeking and collision avoidance. However, several limitations exist. Our current implementation does not support modeling long-term agent intentions. Agents cannot switch actions mid-execution, and action durations are manually defined. Additionally, behaviors are constrained to a predefined list of high-level actions, requiring retraining to accommodate new actions.
For future work, we aim to equip agents with a memory module, ensuring a more natural selection and progression of actions. We would also like to combine Gen-C with a crowd simulation framework, allowing high-level behaviors to control low-level navigation policies, creating more dynamic and realistic simulations. Additionally, we plan to explore combining latent spaces from different datasets, such as "indoor" and "outdoor" scenarios, using techniques such as distillation to enable lifelong learning and dynamic simulation adaptability. We envision that these will allow us to populate virtual worlds with diverse human-like characters, paving the road towards automatic generation of digital crowds that can be applied to a variety of important areas.